OpenStudy (anonymous):

im stuck.someone help? A cone has a base radius of 2 millimeters and a height of 24 millimeters. What is the volume of the cone? A. 32π mm3 B. 384π mm3 C. 96π mm3 D. 16π mm3

OpenStudy (mathmale):

Please type in the formula for the volume of a cone. If unsure, look up this formula (Google "volume of cone."

OpenStudy (anonymous):

v=1/3BH @mathmale

OpenStudy (mathmale):

that'd be fine if you'd please enclose that "1/3" inside parentheses.

OpenStudy (anonymous):

sorry

OpenStudy (mathmale):

What is the formula for the area of a circle (which is what the base is)?

OpenStudy (mathmale):

\[V=\frac{ 1 }{ 3} b h\]

OpenStudy (anonymous):

a=pie(r^2)

OpenStudy (mathmale):

becomes\[V=\frac{ 1 }{ 3 } \pi r^2 h\] both r and h are given, and are in mm. Substitute the given values into this formula (immediately above).

OpenStudy (mathmale):

Yes, the area of the base is \[A _{base}=\pi r^2\]

OpenStudy (anonymous):

And also think about what each variable in your volume formula stands for... B = Area of Base, H = length of Height.

OpenStudy (mathmale):

... and you can see that in my volume formula (above).

OpenStudy (anonymous):

ok..

OpenStudy (mathmale):

\[A _{cone}=\frac{ 1 }{ 3 } \pi (?)^2 (?)\]

OpenStudy (anonymous):

ohhhhhh ok, thanks!:)